OpenStudy (ac3):

Find the area of the region between y=e^X; y=x^(2)+1;x=-1;x=1

OpenStudy (solomonzelman):

maybe y=x²-1 ? because with y=x²+1 u get two different regions

OpenStudy (dumbcow):

 from -1<x<0 , x^2+1 > e^x \[Area = \int\limits_{-1}^0 (x^2 +1) - e^x\] from 0<x<1, e^x > x^2+1 \[Area = \int\limits_0^1 e^x - (x^2 +1)\]